an online store sells two types of speaker docks for smartphones. the higher-priced speaker dock sells for $1900, and the lower-priced speaker dock sells for $800. last week the store sold three times as many lower-priced speaker docks as higher-priced speaker docks. combined sales totaled $34400. indicate how man lower-priced speaker docks did it sell?
Found 2 solutions by Theo, mananth:
Answer by Theo(13342)

higher priced speaker dock is 1900.
lower price speaker dock is 800.
last week store sold 3 times as many lower priced speaker docks as higher priced speaker docks.
let x = number of lower price speaker docks.
let 3x = number of higher price speaker docks.
you get 3x * 800 + x * 1900 = 34400.
combiine like terms to get 2400x + 1900x = 34400.
combine like terms to get 4300x = 34400.
solve for x to get x = 8.
they sold 24 lower price speaker docks and 8 higher price speaker docks.
total sales is 24*800 + 8*1900 = 19200 + 15200 = 34400.
solution is they sold 24 lower priced docks.

Answer by mananth(16946)

the higher-priced speaker dock sells for $1900,
and the lower-priced speaker dock sells for $800.
Let number of higher priced dock be x
Number of lower priced will dock be 3x
Total docks sold = 4x
Total price of both docks = 1000+800 = 1800
1900x +800*(3x) = 34400
1900x +2400x = 34400
4300x = 34400
x= 34400/4300
x= 8 high priced docks
3x= 24 low priced docks
Check 8*1900+24*800 =34400